Could someone please help me with this vectors problem:
Given: a= <-3,4> , b= <2,5>, c= <3,-8>
What is (a+b)-c?
Thank you in advanced.

Answer by jim_thompson5910(35256)

a + b = <-3,4> + <2,5>


a + b = <-3+2,4+5>


a + b = <-1,9>


-------------------------------------------------------


(a+b) - c = <-1,9> - <3,-8>


(a+b) - c = <-1-3,9-(-8)>


(a+b) - c = <-1-3,9+8>


(a+b) - c = <-4,17>